If a plane travels 1600 Km in 1 hour and 12 minutes, how fast is it going in terms of Km/H??
Answer by jim_thompson5910(35256)

12 min = 12/60 hours

12 min = 1/5 hours

12 min = 0.2 hours

---------------------------------

1 hour 12 min = 1 hr + 12 min = 1 hr + 0.2 hr = 1.2 hours


d = rt

1600 = r(1.2)

1600/1.2 = r

1333.333 = r

r = 1333.333

So the speed is roughly 1333.333 km/h